OpenStudy (anonymous):

1/x + 1/a = 1/b; solve for x where a does not equal to b

OpenStudy (anonymous):

multuply both sides by x,a,b

OpenStudy (anonymous):

the answer choices are: a) 1/b - 1/a b) b-a c) 1/ab d) a-b/ab e) ab/a-b

OpenStudy (anonymous):

last one part e

OpenStudy (anonymous):

i don't get how they arrived at the demoninator

OpenStudy (anonymous):

1/x = 1/b - 1/a 1/x = (a-b)/ab x = ab/(a-b)

OpenStudy (anonymous):

after multiplying equation becomes ab+bx=ax ab=ax-bx ab=(a-b)x or multiplying both sides by a-b ab/(a-b)=x